# Decision-Focused Surrogate Modeling for Mixed-Integer Linear Optimization

### v2

@@ 1 Introduction
Effectively operating complex systems in online environments, where input parameters are constantly changing, requires efficient decision making in real time. Many online decision-making frameworks involve the solving of mathematical optimization problems; however, often the computational complexity of the given optimization problem presents a major challenge such that the long solution time renders it ineffective in real-time applications. A common approach to tackling this challenge is to perform the online optimization with a surrogate model, which is an approximation of the original model that can be solved more efficiently (Bhosekar & Ierapetritou, 2018).
@@
Recently, Gupta & Zhang (2024) have developed a new surrogate modeling framework that explicitly aims to construct surrogate models that minimize the decision prediction error defined as the difference between the optimal solutions of the original and the surrogate optimization problems; it is hence referred to as decision-focused surrogate modeling (DFSM). DFSM has been applied to construct convex decision-focused surrogate optimization models (DFSOMs) for nonconvex nonlinear programs that are difficult to solve to global optimality Gupta & Zhang (2024); however, these problems only involve continuous variables. In this work, we extend the DFSM framework to MILPs, which arise in many real-time decision-making problems such as online production scheduling and hybrid model predictive control. Due to their combinatorial nature, MILPs can take long times to solve. Here, we propose to construct DFSOMs in the form of LPs that are computationally significantly more efficient and achieve optimal solutions close to the ones of the original MILPs.
@@ 2.1 ML-based optimization proxies
The line of research that is most closely related to this work is the one on constructing optimization proxies using machine learning (ML). Here, the goal is to learn a direct mapping of the input parameters of an optimization problem to its optimal solution. In that sense, it has the same objective as DFSM, namely to obtain a model that predicts optimal solutions close to the true ones. It uses the same offline generated data consisting of different model inputs and the corresponding true optimal solutions. The main difference is that while DFSM trains a model in the form of a surrogate optimization problem, optimization proxies are ML models, most commonly neural networks (NNs) (Sun et al., 2018; Pan et al., 2019; Karg & Lucia, 2020; Kumar et al., 2021).
Optimization proxies take advantage of the expressive power of deep NNs; however, in deep learning, it is difficult to enforce constraints on the predictions, which often leads to predicted solutions that are infeasible in the original optimization problem. In their work on constructing NN-based optimization proxies for the AC optimal power flow (OPF) problem, Zamzam & Baker (2020) ensure feasibility by generating a training set of strictly feasible solutions through a modified AC OPF formulation and by using the natural bounds of the sigmoid activation function to enforce generation and voltage limits. When also addressing the AC OPF problem, Fioretto et al. (2020) consider constraints in their deep learning framework by augmenting the loss function with penalty terms derived from the Lagrangian dual of the original optimization problem and applying a subgradient method to update the Lagrange multipliers. This approach has also been applied to the security-constrained OPF problem (Velloso & Van Hentenryck, 2021) and job shop scheduling (Kotary et al., 2022). Most of these and similar approaches increase the likelihood of obtaining feasible predictions but cannot guarantee it; hence, in most cases, a feasibility restoration step is added to obtain a feasible solution. One common approach to correcting an infeasible prediction is to project it onto a suitable set such that the projection is a feasible solution (Chen et al., 2018; Paulson & Mesbah, 2020).
@@ 2.2 Decision-focused learning
We now describe a framework that is closely related to the DFSM approach but is not motivated by the need for fast online optimization. Instead, it addresses the following problem: In traditional data-driven optimization, we often follow a two-stage predict-then-optimize approach, i.e. we first predict unknown input parameters from data with external features and then solve the optimization problem with those predicted inputs. Here, the learning step focuses on minimizing the parameter estimation error; however, this does not necessarily lead to the best decisions (evaluated with the true parameter values) in the optimization step. In contrast, decision-focused learning (Wilder et al., 2019), also known as smart predict-then-optimize (Elmachtoub & Grigas, 2022) or predict-and-optimize (Mandi et al., 2020), integrates the two steps to explicitly account for the quality of the optimization solution in the learning of the model parameters.
Elmachtoub & Grigas (2022) consider decision-focused learning for optimization problems with a linear objective function and a closed convex feasible region. They develop a convex surrogate loss function that is Fisher consistent with respect to the original loss function and allows the use of stochastic gradient descent to efficiently solve the problem. The same algorithm has also been applied to combinatorial problems (Mandi et al., 2020). Decision-focused learning has further motivated research on differentiable optimization in deep learning (Amos & Kolter, 2017), which allows NNs to comprise differentiable layers that represent full optimization problems. Using this approach, quadratic programs (Donti et al., 2017), linear programs (Wilder et al., 2019), and mixed-integer linear programs (Ferber et al., 2020) have been considered in the training of decision-focused NNs. This approach is applied by Ferber et al. (2023) to address a problem of similar nature as our DFSM problem, namely to learn linear surrogate objective functions for discrete optimization problems with linear constraints but nonlinear objective functions.
@@ 2.3 ML-enhanced mixed-integer optimization
There is a rapidly growing body of literature on using ML to speed up the solution of mixed-integer optimization problems (Bengio et al., 2021). Our work as well as the ML-based optimization proxies reviewed in Section 2.1 fall into this broad category. Other strategies include using ML to learn more effective primal heuristics (Bengio et al., 2020; Shen et al., 2021), branching rules (Lodi & Zarpellon, 2017; Khalil et al., 2016), the set of active constraints at the optimal solution (Bertsimas & Stellato, 2022), and how to warm-start MILPs (Jiménez-Cordero et al., 2022). More closely related to our work are the contributions on learning how to improve the generation of cutting planes (Deza & Khalil, 2023; Tang et al., 2020; Huang et al., 2022). However, while these methods add cuts as part of a branch-and-cut algorithm, our approach learns parametric cuts that can be added immediately to the LP relaxation for a given model input, as discussed in more detail in Section 3.
@@ 2.4 Multiparametric programming
While DFSM and optimization proxies can generally only provide approximate optimal solutions, multiparametric programming is an exact approach that maps model inputs directly to the true optimal solutions through explicit functions (Oberdieck et al., 2016). The solution of a multiparametric LP follows from the basic sensitivity theorem (Fiacco & Ishizuka, 1990), which states that the active set remains unchanged in the neighborhood of the given vector of input parameters and the corresponding optimal solution. This results in the partitioning of the feasible parameter space into polytopic regions called critical regions. The goal of multiparametric programming is to obtain these critical regions and their corresponding optimal policies offline, then the online optimization reduces to selecting and applying the right policy for a given parameter vector that lies in the corresponding critical region.
When solving multiparametric MILPs, the same concept can be applied. One naive way is to enumerate all the possible discrete solutions and then treat the multiparametric MILP as a set of many multiparametric LPs (Roodman, 1972), one for each discrete solution. Thus, the resulting solution to the multiparametric MILP is again achieved by partitioning the parameter space into critical regions, each of which is associated with an optimal policy as an affine function of the model parameters. More advanced methods apply algorithms based on branch-and-bound (Ohtake & Nishida, 1985; Acevedo & Pistikopoulos, 1999) and decomposition (Dua & Pistikopoulos, 2000).
Multiparametric programming has found several successful applications in explicit model predictive control (Alessio & Bemporad, 2009; Pistikopoulos et al., 2015) with extensions to other frameworks such as integrated design and operations (Burnak et al., 2019) and optimization under uncertainty (Charitopoulos et al., 2018). However, as the problem size grows with the number of decision variables, constraints, and model parameters, the construction of the full multiparametric solution becomes challenging as the number of critical regions grows exponentially with the problem size. Moreover, given a parameter vector, finding the critical region for that point becomes a point location problem, which can also be a computationally difficult problem for larger systems with many critical regions.
@@ 3.1 General formulation
where =$\mathcal{I}=\{(\bar{u}_{i},x^{*}_{i})\}_{i=1}^{N}$ constitutes the training dataset with $x^{*}_{i}$ being the optimal solution to the original MILP (1) for input $\bar{u}_{i}$. As shown in (3a), we use the $\ell_{1}$-norm to measure the decision error, and the objective is to choose $\theta$ from some appropriate set $\Theta$ such that the sum of decision errors across all training data points is minimized. Problem (3) is a bilevel optimization problem with $N$ lower-level problems. As shown in (3b), each lower-level problem represents the surrogate LP for a particular input $\bar{u}_{i}$, and $\tilde{x}_{i}$ is constrained to be an optimal solution to that LP.
@@ 3.2 Solution strategy
Problem (4) is a nonconvex NLP that violates constraint qualification (due to, for example, the complementary slackness conditions), which means that standard NLP solvers cannot guarantee convergence to the optimal solution when directly applied to this problem. Here, we adopt the solution strategy proposed by Gupta & Zhang (2022) for data-driven inverse optimization problems, which turn out to have the same structure as the DFSM problem. We apply an exact penalty reformulation that is achieved by penalizing the violation of constraints (4b)-(4f) in the objective function using the $\ell_{1}$-norm. We can then iteratively update the penalty parameter until no more constraint violation is observed at which point the algorithm is guaranteed to have converged to a local solution of problem (4). The penalty reformulation takes the form of a multiconvex optimization problem in terms of the three sets of variables $\theta$, $\tilde{x}$, and $(\lambda,\mu)$, i.e. the problem is convex with respect to each of these three sets of variables. This structure can be exploited using a block coordinate descent (BCD) algorithm that allows an efficient decomposition of the problem, which has been crucial in solving large instances of the DFSM problem. For more details on the solution algorithm, see Appendix A.2.
@@ 4.1.1 Data generation and surrogate design
We also compare the performance of the DFSM approach with that of three NN-based optimization proxies, one that applies a simple feedforward NN architecture, a second one where the loss function contains penalty terms derived from the Lagrangian dual of the optimization problem to impose feasibility on the predicted solutions (Fioretto et al., 2020), and a third one that is based on a graph NN (GNN) architecture. To represent the MILP as a GNN, we derive a bipartite graph whose two sets of nodes correspond to the problem’s variables and constraints, respectively. An edge connects a variable node with a constraint node if that variable appears in the constraint. Also, features such as variable bounds and right-hand-side values are assigned to the nodes to fully describe the MILP. These NN-based optimization proxies are similarly trained in a decision-focused manner using the same datasets. See Appendix A.3 for more details on the design and training of the NN-based optimization proxies.
@@ A.1 Illustrative example
$\underset{x_{1}\in\mathbb{Z},\,x_{2}\in\mathbb{Z}}{\mathrm{maximize}}$ $4.8x_{1}+6x_{2}$ (7)
$\mathrm{subject\;to}$ $4x_{1}+3x_{2}\leq 70$
$680\displaystyle 100ux_{1}+85x_{2}\leq 800u+$ 680
$x_{1}\leq 17,\;x_{2}\leq 17.$
@@
$\underset{x_{1}\in\mathbb{R},\,x_{2}\in\mathbb{R}}{\mathrm{maximize}}$ $4.8x_{1}+6x_{2}$ (8)
$\mathrm{subject\;to}$ $4x_{1}+3x_{2}\leq 70$
$680\displaystyle 100ux_{1}+85x_{2}\leq 800u+$ 680
$(17.9-56.9{u}+45.1{u}^{2})x_{1}+(-4.5+25.3{u}-14.3u^{2})x_{2}\leq 100$
$(7.4+6.4{u}-7.9{u}^{2})x_{1}\leq 100$
$(-0.2+14.5{u}-9.4{u}^{2})x_{1}+(12-9.7{u}+3.7u^{2})x_{2}\leq 100$
$x_{1}\leq 17,\;x_{2}\leq 17.$
@@ A.3.1 Feedforward NN architecture
For the feedforward NN, the numbers of hidden layers and neurons in each layer are determined through a trial-and-error approach aiming to minimize the training loss over a fixed number of epochs. The number of hidden layers is based on the empirical observation that increasing the number of hidden layers resulted in little improvements in the model’s accuracy. The design of the specific NN model used in the hybrid vehicle control case study is as follows: the input layer contains $T$ nodes, the same as the dimension of the input vector $u$. The two hidden layers have $3T$ and $6T$ nodes, and finally, the output layer consists of $13T+$ 1 nodes aligning with the dimensionality of the decision space. We also account for the effect of different learning rates (namely 0.1, 0.01, and 0.001) on the training loss. Based on our analysis, we determine 0.01 to be the most suitable learning rate for our purpose.
@@ A.3.3 GNN architecture
    The features of each constraint node contain information about the corresponding right-hand-side value and whether the constraint is an inequality ()($\leq$)( ≤ ) or an equality (=)(=)( = ).

### v3

@@ 1 Introduction
Effectively operating complex systems in online environments, where input parameters are constantly changing, requires efficient decision making in real time. Many online decision-making frameworks involve the solving of mathematical optimization problems; however, often the computational complexity of the given optimization problem presents a major challenge such that the long solution time renders it ineffective in real-time applications. A common approach to tackling this challenge is to perform the online optimization with a surrogate model, which is an approximation of the original model that can be solved more efficiently (Bhosekar2018).
@@
Recently, Gupta and Zhang (2024) have developed a new surrogate modeling framework that explicitly aims to construct surrogate models that minimize the decision prediction error defined as the difference between the optimal solutions of the original and the surrogate optimization problems; it is hence referred to as decision-focused surrogate modeling (DFSM). DFSM has been applied to construct convex decision-focused surrogate optimization models (DFSOMs) for nonconvex nonlinear programs that are difficult to solve to global optimality Gupta and Zhang (2024); however, these problems only involve continuous variables. In this work, we extend the DFSM framework to MILPs, which arise in many real-time decision-making problems such as online production scheduling and hybrid model predictive control. Due to their combinatorial nature, MILPs can take long times to solve. Here, we propose to construct DFSOMs in the form of LPs that are computationally significantly more efficient and achieve optimal solutions close to the ones of the original MILPs.
@@ 2.1 ML-based optimization proxies
The line of research that is most closely related to this work is the one on constructing optimization proxies using machine learning (ML). Here, the goal is to learn a direct mapping of the input parameters of an optimization problem to its optimal solution. In that sense, it has the same objective as DFSM, namely to obtain a model that predicts optimal solutions close to the true ones. It uses the same offline generated data consisting of different model inputs and the corresponding true optimal solutions. The main difference is that while DFSM trains a model in the form of a surrogate optimization problem, optimization proxies are ML models, most commonly neural networks (NNs) (Sun2018; Pan2019; Karg2020; Kumar2021).
Optimization proxies take advantage of the expressive power of deep NNs; however, in deep learning, it is difficult to enforce constraints on the predictions, which often leads to predicted solutions that are infeasible in the original optimization problem. In their work on constructing NN-based optimization proxies for the AC optimal power flow (OPF) problem, Zamzam2020 ensure feasibility by generating a training set of strictly feasible solutions through a modified AC OPF formulation and by using the natural bounds of the sigmoid activation function to enforce generation and voltage limits. When also addressing the AC OPF problem, Fioretto2020 consider constraints in their deep learning framework by augmenting the loss function with penalty terms derived from the Lagrangian dual of the original optimization problem and applying a subgradient method to update the Lagrange multipliers. This approach has also been applied to the security-constrained OPF problem (Velloso2021) and job shop scheduling (Kotary2022). Most of these and similar approaches increase the likelihood of obtaining feasible predictions but cannot guarantee it; hence, in most cases, a feasibility restoration step is added to obtain a feasible solution. One common approach to correcting an infeasible prediction is to project it onto a suitable set such that the projection is a feasible solution (Chen2018a; Paulson2020).
@@ 2.2 Decision-focused learning
We now describe a framework that is closely related to the DFSM approach but is not motivated by the need for fast online optimization. Instead, it addresses the following problem: In traditional data-driven optimization, we often follow a two-stage predict-then-optimize approach, i.e. we first predict unknown input parameters from data with external features and then solve the optimization problem with those predicted inputs. Here, the learning step focuses on minimizing the parameter estimation error; however, this does not necessarily lead to the best decisions (evaluated with the true parameter values) in the optimization step. In contrast, decision-focused learning (Wilder2019), also known as smart predict-then-optimize (Elmachtoub2022) or predict-and-optimize (Mandi2020), integrates the two steps to explicitly account for the quality of the optimization solution in the learning of the model parameters.
Elmachtoub2022 consider decision-focused learning for optimization problems with a linear objective function and a closed convex feasible region. They develop a convex surrogate loss function that is Fisher consistent with respect to the original loss function and allows the use of stochastic gradient descent to efficiently solve the problem. The same algorithm has also been applied to combinatorial problems (Mandi2020). Decision-focused learning has further motivated research on differentiable optimization in deep learning (Amos2017), which allows NNs to comprise differentiable layers that represent full optimization problems. Using this approach, quadratic programs (Donti2017), linear programs (Wilder2019), and mixed-integer linear programs (Ferber2020) have been considered in the training of decision-focused NNs. This approach is applied by Ferber et al. (2023) to address a problem of similar nature as our DFSM problem, namely to learn linear surrogate objective functions for discrete optimization problems with linear constraints but nonlinear objective functions.
@@ 2.3 ML-enhanced mixed-integer optimization
There is a rapidly growing body of literature on using ML to speed up the solution of mixed-integer optimization problems (Bengio et al., 2021). Our work as well as the ML-based optimization proxies reviewed in Section 2.1 fall into this broad category. Other strategies include using ML to learn more effective primal heuristics (Bengio et al., 2020; Shen et al., 2021), branching rules (Lodi and Zarpellon, 2017; Khalil et al., 2016), the set of active constraints at the optimal solution (Bertsimas and Stellato, 2022), and how to warm-start MILPs (Jiménez-Cordero et al., 2022). More closely related to our work are the contributions on learning how to improve the generation of cutting planes (Deza and Khalil, 2023; Tang et al., 2020; Huang et al., 2022). However, while these methods add cuts as part of a branch-and-cut algorithm, our approach learns parametric cuts that can be added immediately to the LP relaxation for a given model input, as discussed in more detail in Section 3.
@@ 2.4 Multiparametric programming
While DFSM and optimization proxies can generally only provide approximate optimal solutions, multiparametric programming is an exact approach that maps model inputs directly to the true optimal solutions through explicit functions (Oberdieck2016). The solution of a multiparametric LP follows from the basic sensitivity theorem (Fiacco and Ishizuka, 1990), which states that the active set remains unchanged in the neighborhood of the given vector of input parameters and the corresponding optimal solution. This results in the partitioning of the feasible parameter space into polytopic regions called critical regions. The goal of multiparametric programming is to obtain these critical regions and their corresponding optimal policies offline, then the online optimization reduces to selecting and applying the right policy for a given parameter vector that lies in the corresponding critical region.
When solving multiparametric MILPs, the same concept can be applied. One naive way is to enumerate all the possible discrete solutions and then treat the multiparametric MILP as a set of many multiparametric LPs (Roodman, 1972), one for each discrete solution. Thus, the resulting solution to the multiparametric MILP is again achieved by partitioning the parameter space into critical regions, each of which is associated with an optimal policy as an affine function of the model parameters. More advanced methods apply algorithms based on branch-and-bound (Ohtake and Nishida, 1985; Acevedo and Pistikopoulos, 1999) and decomposition (Dua and Pistikopoulos, 2000).
Multiparametric programming has found several successful applications in explicit model predictive control (Alessio2009; Pistikopoulos2015) with extensions to other frameworks such as integrated design and operations (Burnak et al., 2019) and optimization under uncertainty (Charitopoulos et al., 2018). However, as the problem size grows with the number of decision variables, constraints, and model parameters, the construction of the full multiparametric solution becomes challenging as the number of critical regions grows exponentially with the problem size. Moreover, given a parameter vector, finding the critical region for that point becomes a point location problem, which can also be a computationally difficult problem for larger systems with many critical regions.
@@ 3.1 General formulation
where $\mathcal{I}=\{(\bar{u}_{i},x^{*}_{i})\}_{i=1}^{N}$ constitutes the training dataset with $x^{*}_{i}$ being the optimal solution to the original MILP (1) for input $\bar{u}_{i}$. As shown in (3a), we use the $\ell_{1}$-norm to measure the decision error, and the objective is to choose $\theta$ from some appropriate set $\Theta$ such that the sum of decision errors across all training data points is minimized. Problem (3) is a bilevel optimization problem with $N$ lower-level problems. As shown in (3b), each lower-level problem represents the surrogate LP for a particular input $\bar{u}_{i}$, and $\tilde{x}_{i}$ is constrained to be an optimal solution to that LP.
@@ 3.2 Solution strategy
Problem (4) is a nonconvex NLP that violates constraint qualification (due to, for example, the complementary slackness conditions), which means that standard NLP solvers cannot guarantee convergence to the optimal solution when directly applied to this problem. Here, we adopt the solution strategy proposed by Gupta and Zhang (2022) for data-driven inverse optimization problems, which turn out to have the same structure as the DFSM problem. We apply an exact penalty reformulation that is achieved by penalizing the violation of constraints (4b)-(4f) in the objective function using the $\ell_{1}$-norm. We can then iteratively update the penalty parameter until no more constraint violation is observed at which point the algorithm is guaranteed to have converged to a local solution of problem (4). The penalty reformulation takes the form of a multiconvex optimization problem in terms of the three sets of variables $\theta$, $\tilde{x}$, and $(\lambda,\mu)$, i.e. the problem is convex with respect to each of these three sets of variables. This structure can be exploited using a block coordinate descent (BCD) algorithm that allows an efficient decomposition of the problem, which has been crucial in solving large instances of the DFSM problem. For more details on the solution algorithm, see Appendix A.2.
@@ 4.1.1 Data generation and surrogate design
We also compare the performance of the DFSM approach with that of three NN-based optimization proxies, one that applies a simple feedforward NN architecture, a second one where the loss function contains penalty terms derived from the Lagrangian dual of the optimization problem to impose feasibility on the predicted solutions (Fioretto2020), and a third one that is based on a graph NN (GNN) architecture. To represent the MILP as a GNN, we derive a bipartite graph whose two sets of nodes correspond to the problem’s variables and constraints, respectively. An edge connects a variable node with a constraint node if that variable appears in the constraint. Also, features such as variable bounds and right-hand-side values are assigned to the nodes to fully describe the MILP. These NN-based optimization proxies are similarly trained in a decision-focused manner using the same datasets. See Appendix A.3 for more details on the design and training of the NN-based optimization proxies.
@@ A.1 Illustrative example
[…]
@@
[…]
@@ A.3.1 Feedforward NN architecture
For the feedforward NN, the numbers of hidden layers and neurons in each layer are determined through a trial-and-error approach aiming to minimize the training loss over a fixed number of epochs. The number of hidden layers is based on the empirical observation that increasing the number of hidden layers resulted in little improvements in the model’s accuracy. The design of the specific NN model used in the hybrid vehicle control case study is as follows: the input layer contains $T$ nodes, the same as the dimension of the input vector $u$. The two hidden layers have $3T$ and $6T$ nodes, and finally, the output layer consists of $3T+1$ nodes aligning with the dimensionality of the decision space. We also account for the effect of different learning rates (namely 0.1, 0.01, and 0.001) on the training loss. Based on our analysis, we determine 0.01 to be the most suitable learning rate for our purpose.
@@ A.3.3 GNN architecture
    The features of each constraint node contain information about the corresponding right-hand-side value and whether the constraint is an inequality $(\leq)$ or an equality $(=)$.